# Second-Generation Planet Formation in Post-AGB Discs: Testing the Role of Gravitational Instability

### v1

@@ 7 Conclusion
    Under realistic assumptions of disc mass 0.020.050.02-0.05M, disc structure, and density and temperature profiles, post-AGB discs are generally gravitationally stable at present. Even systems with the most optimistic parameters—such as IRAS 08544-4431 with a disc mass of 0.10.1M—remain above the fragmentation threshold, with Toomre $Q>1$ throughout the disc, largely due to thermal support from the luminous central star. In contrast, discs around lower-luminosity central objects (e.g. white dwarfs or Sun-like stars) can reach the gravitationally unstable regime at lower masses.
@@
    Using self-similar viscous disc evolution models, in which the surface density profile retains its shape while the overall scale (mass and characteristic radius) evolves due to viscous spreading, we explored whether post-AGB discs may have been unstable earlier in their lifetimes. We found that, even after evolving them backward in time across several viscous timescales, most remain stable, unlike some protoplanetary discs (e.g. HL Tauri, Elias 2-27, GQ Lupi), which likely passed through an unstable phase during earlier epochs. Our analysis also shows that higher viscosity parameters ($\alpha\sim 10^{-2}$) are more consistent with expected post-AGB disc lifetimes.
@@
    Revisiting the post-common envelope binary NN Serpentis, we confirm that if the disc parameters proposed by 2014schleicher are assumed, the disc would become gravitationally unstable beyond  $\sim\,3$ au. However, the planet masses predicted by their gravitationally induced clump growth model slightly underpredict those of the observed planet candidates, and recent developments in common envelope theory suggest that fallback discs may be far less massive than previously assumed.
It is important to note that using the Toomre parameter as a diagnostic for gravitational instability, while useful, does not take factors such as vertical structure, turbulence, and binary-driven dynamics into account. Post-AGB disks can differ from YSO disks in these regards as well, hence the results of this study should be treated as a first order analysis.
Overall, our analysis demonstrates that gravitational instability is highly unlikely to account for second-generation planet formation in typical post-AGB discs, as fragmentation would require unrealistically high disc masses ($\gtrsim 0.4$ M), even under compact and cool conditions. The observed signatures of planets or planet-like companions in several post-AGB systems therefore point towards alternative formation channels—most likely a modified version of core accretion operating on shorter timescales in evolved environments. This study provides a first theoretical baseline for interpreting such systems. While future observations will better constrain disc masses, gas-to-dust ratios, temperatures, sizes, and lifetimes—and may uncover additional evidence for planet formation—we will, in parallel, explore additional theoretical models to investigate planet formation pathways in post-main-sequence environments.
@@ Acknowledgments
AP acknowledges the financial support provided by the International Macquarie Research Excellence Scholarship (iMQRES) program, received throughout the duration of this research. DK, ODM and MW acknowledge the support from the Australian Research Council Discovery Project DP240101150.
@@ Appendix A Parameters of YSO systems
In this section we present the studies we have used to compile the YSO disc properties into Table 1.
We note that the reported inner radii in the literature typically refer to the inner radius of the dust component. Some studies explicitly report an inner gas cavity (such as HD169142), but others do not distinguish between the two. If a study has not stated that the inner radius they have reported is the dust radius or gas radius, we have taken that as the gas inner radius. In the case of some studies that carried out simulations and had reported the inner radius they had taken for the disk cavity, such as AB Aurigae, and Elias 2-27, we used the value they adopted. For PDS 70 there were no inner cavities in literature, hence we adopted $r_{\rm in}=0$.
There are two different reported masses for the disc of AB Aurigae, each obtained with different methods, and we have carried out Toomre analysis with both values. rivier2024 report $3.2\times 10^{-3}~\rm M_{\odot}$ which is based on ALMA and VLA observations to find the dust mass and then use a gas-to-dust ratio to estimate the total disc mass. On the other hand, 2023Speedie have estimated the disc mass based on their observation and subsequent modelling of the gravitational wiggle they observe in the disc, which they argue is a signature of gravitational instability, and they find a disc mass of $0.7~\rm M_{\odot}$ best fits their model. Other than the disc mass, the rest of the tabulated values are those that 2023Speedie have taken for their hydrodynamics simulation (note that this is the only system in this table which uses a power law for surface density rather than Equation 13).
For HL Tauri, the parameters for the star’s radius and luminosity are from 1999menshchikov. The star mass was calculated in 2016pinte by fitting models to CO and HCO+ velocities obtained from ALMA, and the rest of the parameters, most notably the disc mass, were taken from Booth_2020, who have used $^{13}\rm C^{17}\rm O$ measurements from ALMA to model and find the disc mass which is significantly more massive compared to previous measurements.
For Elias 2-27 , we adopted the value of the inner radius of the disc from Paneque_Carre_o_2021, the photospheric temperature and radius of the star from 2009Andrews and the rest from Veronesi_2021, who used ALMA observations to find the disc and star masses. It is noteworthy to mention that this system has been also analyzed to see whether it is prone to instabilities in Perez_2016, albeit with different values, and for their standard opacity value their resulting Toomre estimates are similar to us, noting the slightly different values they have adopted for the disc parameters.
@@
For HD 169142, we take the values of star mass, radius, and temperature from Quanz_2013. Quanz_2013 report 3 different temperatures, and the adopted one was calculated from high-resolution spectroscopy by 2006guimaraes who find $T_{\rm eff}=7500$K. We take the characteristic disc radius and other disc parameters from Fedele_2017, who used a best fit to ALMA observations to estimate the disc parameters. For the disc mass, Fedele_2017 mentioned that the total dust mass is $10^{-4}$ M and the best fit they find for gas-to-dust mass ratio is 80, which gives the disc mass to be 0.008 M. They also determined $\Sigma_{\rm ch}=6.5$ g/cm2 as the surface density at the characteristic radius.
For GQ Lup, both disc and star parameters had been summarized by Wu_2017. The disc surface density was fitted using ALMA data. Similarly for L1448 IRS3B, the disc and star parameters have been taken from Reynolds_2021, who used ALMA data.
Appendix B Planet Formation Around Post-Common Envelope Binary Systems
For completeness and consistency we calculate here the mass and orbital parameters of the two planets using a realistic temperature profile that allows us not to assume a value of $Q$, and compare our values with those of 2014schleicher.
2014schleicher started their analysis of the NN Ser post-common envelope (CE) binary and its putative planets, by motivating their choice of disk mass ($\sim 0.1$ M). They then estimated the amount of angular momentum deposited by the companion into the envelope of the AGB star, $L_{\rm dep}$:
@@ Appendix B Planet Formation Around Post-Common Envelope Binary Systems
where $M_{1}$ and $R_{*}$ are the AGB star’s mass and radius, respectively and $M_{2}$ is the companion’s mass (see Table 3). In order to account for additional angular momentum that the envelope may already have had at the time of CE in-spiral, they assumed that the actual disk’s specific angular momentum would be a factor of a few larger than the one deposited by the companion. They parameterised this in the following way:
@@
where $\alpha_{\rm L}$ is the enhancement factor for the specific angular momentum, $L_{\rm disc}$ is the angular momentum of the disc and $M_{\rm ej}$ is the mass ejected during the CE phase, but that presumably remains bound and forms the disk.
After the disc has formed, they assumed that its surface density, $\Sigma$, settles into a power-law radial profile with a truncation radius, $r_{\rm out}$, described by Equation 10. If they calculate the angular momentum of the disc assuming a Keplerian velocity surrounding a stripped core with mass $M_{\rm c}$ and a companion star of mass $M_{2}$, they could find the outer radius, $r_{\rm out}$:
@@
2014schleicher made the assumption that this disk is gravitationally unstable ($Q=1$). From an initial clump mass, $M_{\rm cl}$, approximately estimated as:
@@
with $\alpha_{\rm crit}$ the critical viscosity parameter. This equation is obtained from the gap-opening mass expression of 1986lin [^13] Equation 34 of 2014schleicher has a missing factor of $\sqrt{\pi}$ compared to that of 1986lin.. They obtain masses of 6.6 MJup and 1.9 MJup (cf. observed values of 7 MJup and 1.7 MJup) for orbital separations of 5.4 au and 3.4 au.
@@
We then estimate the final clump mass (Equations 30 and 31) by using a realistic temperature profile and not assuming that $Q=1$ (Equation 7). The predicted final clump masses for orbital separations of 5.4 au and 3.4 au (blue line) are 2 MJup and 1.2 MJup, somewhat lower than the values obtained by 2014schleicher. Despite this small discrepancy, the analysis of 2014schleicher is consistent with what we have found so that we can conclude that if the disk mass is close to their adopted 0.1 M then indeed this disk could make the observed planets by gravitational fragmentation.
Figure 11: Clump masses and final (planet) masses as a function of orbital distance for NN Serpentis using the values used by 2014schleicher, but assuming the temperature profile of Equation 7 and $r_{\rm out}=7.5$ au.

### v2

@@ 7 Conclusion
    Under realistic assumptions of disc mass $\sim 0.02$ M, disc structure, and density and temperature profiles, post-AGB discs are generally gravitationally stable at present. Even systems with the most optimistic parameters - such as IRAS 08544-4431 with a disc mass of 0.10.1M - remain above the fragmentation threshold, with Toomre $Q>1$ throughout the disc, largely due to thermal support from the luminous central star. In contrast, discs around lower-luminosity central objects (e.g., white dwarfs or Sun-like stars) can reach the gravitationally unstable regime at lower masses.
@@
    Using self-similar viscous disc evolution models, in which the surface density profile retains its shape while the overall scale (mass and characteristic radius) evolves due to viscous spreading, we explored whether post-AGB discs may have been unstable earlier in their lifetimes. We found that, even after evolving them backward in time across several viscous timescales, most remain stable, unlike some protoplanetary discs (e.g., HL Tauri, Elias 2-27, GQ Lupi), which likely passed through an unstable phase during earlier epochs. Our analysis also shows that higher viscosity parameters ($\alpha\sim 10^{-2}$) are more consistent with expected post-AGB disc lifetimes.
@@
    Revisiting the post-common envelope binary NN Ser, we confirm that if the disc parameters proposed by 2014schleicher are assumed, the disc would become gravitationally unstable beyond  $\sim\,3$ au. However recent developments in common envelope theory suggest that fallback discs may be far less massive than previously assumed.
It is important to note that using the Toomre parameter as a diagnostic for gravitational instability, while useful, does not take factors such as vertical structure, turbulence, and binary-driven dynamics into account. Post-AGB discs can differ from YSO discs in these regards as well, hence the results of this study should be treated as a first order analysis.
Overall, our analysis demonstrates that gravitational instability is highly unlikely to account for second generation planet formation in typical post-AGB discs, as fragmentation would require unrealistically high disc masses ($\gtrsim 0.4$ M), even under compact and cool conditions. The observed signatures of planets or planet-like companions in several post-AGB systems therefore point towards alternative formation channels - most likely a modified version of core accretion operating on shorter timescales in evolved environments. This study provides a first theoretical baseline for interpreting such systems. While future observations will better constrain disc masses, gas-to-dust ratios, temperatures, sizes, and lifetimes - and may uncover additional evidence for planet formation - we will, in parallel, explore additional theoretical models to investigate planet formation pathways in post-main-sequence environments.
@@ Acknowledgments
AP acknowledges the financial support provided by the International Macquarie Research Excellence Scholarship (iMQRES) program, received throughout the duration of this research. DK, ODM and MW acknowledge the support from the Australian Research Council Discovery Project DP240101150. We thank Toon De Prins for carefully reading the manuscript and pointing out several typographical errors. We also thank the anonymous referee for valuable comments that improved the paper.
@@ Appendix A Parameters of YSO systems
In this section we present the studies we have used to compile the YSO disc properties in Table 1.
We note that the reported inner radii in the literature typically refer to the inner radius of the dust component. Some studies explicitly report an inner gas cavity (such as HD169142), but others do not distinguish between the two. If a study does not state whether the reported inner radius is the dust or gas radius, we have taken it to be the gas inner radius. For studies that carried out simulations and reported their adopted inner radius, such as for AB Aurigae, or Elias 2-27, we have used their value. For PDS 70 there were no reported inner cavities in the literature, hence we adopted $r_{\rm in}=0$.
There are two different reported masses for the disc of AB Aurigae, each obtained with different methods, and we have carried out Toomre analysis with both values. rivier2024 reported $3.2\times 10^{-3}~\rm M_{\odot}$ which is based on ALMA and VLA observations to find the dust mass and then used a gas-to-dust ratio to estimate the total disc mass. On the other hand, 2023Speedie have estimated the disc mass based on their observation and subsequent modelling of the gravitational wiggle they observe in the disc, which they argued is a signature of gravitational instability, and found that a disc mass of $0.7~\rm M_{\odot}$ best fits their model. Other than the disc mass, the rest of the tabulated values are those that 2023Speedie have taken for their hydrodynamics simulation (note that AB Aurigae and L1448 IRS3B are the only systems which were modelled using a power law for the surface density rather than Equation 13).
For HL Tauri, the parameters for the star’s radius and luminosity are from 1999menshchikov. The star’s mass was calculated by 2016pinte by fitting models to CO and HCO+ velocities obtained from ALMA, and the rest of the parameters, most notably the disc mass, were taken from Booth_2020, who have used $^{13}\rm C^{17}\rm O$ measurements from ALMA to find a disc mass that is significantly more massive compared to previous measurements.
For Elias 2-27 , we adopt the value of the inner radius of the disc from Paneque_Carre_o_2021, the photospheric temperature and radius of the star from 2009Andrews and the rest from Veronesi_2021, who used ALMA observations to find the disc and star masses. It is noteworthy to mention that this system has been also analysed for unstable behaviour by Perez_2016, albeit with different values, and for their standard opacity value their resulting Toomre estimates are similar to ours, noting the slightly different values they have adopted for the disc parameters.
@@
For HD 169142, we take the values of star mass, radius, and temperature from Quanz_2013. Quanz_2013 reported three different stellar surface temperatures, and the adopted one was calculated from high resolution spectroscopy by 2006guimaraes, who found $T_{\rm eff}=7500$ K. We take the characteristic disc radius and other disc parameters from Fedele_2017, who used a best fit to ALMA observations to estimate the disc parameters. For the disc mass, Fedele_2017 mentioned that the total dust mass is $10^{-4}$ M and the best fit they find for the gas-to-dust mass ratio is 80, which results in a disc mass of 0.008 M. They also determined $\Sigma_{\rm ch}=6.5$ g/cm2 to be the surface density at the characteristic radius.
For GQ Lup, both disc and star parameters have been summarised by Wu_2017. The disc surface density was fitted using ALMA data. Similarly for L1448 IRS3B, the disc and star parameters have been taken from Reynolds_2021, who used ALMA data.
Appendix B Benchmarking Against Published Toomre Q Profiles
To further test the robustness of our prescription, we compute the Toomre parameter for three additional discs (DM Tau, LkCa 15, and GM Aur) that have published Q profiles (Martire_2024; Longarini_2025, e.g.,) and published stellar radii and temperatures 2011Andrews, enabling us to reproduce their temperature structures within our framework. The resulting Q profiles are presented in Figure 10 and show a similar overall shape, although our profiles display smaller minima of Q=23Q=2-3 as opposed to their values of Q=78Q=7-8. This relatively minor discrepancy arises from the different temperature profiles adopted by them (individually derived via modelling), as opposed to those we calculate using Equation 7).
Figure 10: Toomre parameter $Q$ versus disc radius for the YSO discs of DM Tau, GM Aur, and LkCa 15. The mass, and characteristic radius of these systems were taken from Martire_2024; Longarini_2025, whilst their star temperature and radius were taken from 2011Andrews. We used the surface density profile of viscously evolving discs (Equation 13) to model these systems, assuming no inner cavities for them. See Figure 7 of Longarini_2025, and Figure 9 of Martire_2024 for comparison.
Appendix C Planet Formation Around Post-Common Envelope Binary Systems
For completeness and consistency we calculate here the mass and orbital parameters of the two planets of NN Ser using a realistic temperature profile that allows us not to assume a value of $Q$, and compare our values with those of 2014schleicher.
2014schleicher started their analysis of the NN Ser post-common envelope (CE) binary and its putative planets, by motivating their choice of disc mass ($\sim 0.1$ M).
They then estimated the amount of angular momentum deposited by the companion into the envelope of the AGB star, $L_{\rm dep}$:
@@ Appendix C Planet Formation Around Post-Common Envelope Binary Systems
where $M_{1}$ and $R_{*}$ are the AGB star’s mass and radius, respectively and $M_{2}$ is the companion’s mass (see Table 3). In order to account for additional angular momentum that the envelope may already have had at the time of CE in-spiral, they assumed that the actual disc’s specific angular momentum would be a factor of a few larger than the one deposited by the companion. They parameterised this in the following way:
@@
where $\alpha_{\rm L}$ is the enhancement factor for the specific angular momentum, $L_{\rm disc}$ is the angular momentum of the disc and $M_{\rm ej}$ is the mass ejected during the CE phase, but that presumably remains bound and forms the disc.
After the disc has formed, they assumed that its surface density, $\Sigma$, settles into a power-law radial profile with a truncation radius, $r_{\rm out}$, described by Equation 10. If they calculate the angular momentum of the disc assuming Keplerian velocity surrounding a stripped core with mass $M_{\rm c}$ and a companion star of mass $M_{2}$, they could find the outer radius, $r_{\rm out}$:
@@
Table 3: Parameters adopted or determined by 2014schleicher in their study of NN Ser.
[…]
Disc Mass, $M_{\rm disk}$ (M) 0.146
Ejected Mass, $M_{\rm ej}$ (M ) 1.24
Primary mass, $M_{1}$ (M  ) 2
Companion mass, $M_{2}$ (M ) 0.111
Primary core mass, $M_{\rm c}$ (M  ) 0.535
Primary radius, $R_{*}$ (R) 185
Disc’s surface density index, $n$ 1
Disc’s outer radius, $r_{\rm out}$ (R) 7.5  [^12] The value we find for $r_{\rm out}$ and $L_{\rm dep}$ is different from that of 2014schleicher, see text for more details.
Disc’s ang. mom., $L_{\rm dep}$ (g cm s-1) $1.2\times 10^{52}$ 12
Disc’s ang. mom. enhancement factor, $\alpha_{\rm L}$ 12.5
critical viscosity, $\alpha_{\rm crit}$ 0.3
2014schleicher made the assumption that this disc is gravitationally unstable ($Q=1$). From an initial clump mass, $M_{\rm cl}$, approximately estimated as:
@@
with $\alpha_{\rm crit}$ the critical viscosity parameter. This equation is obtained from the gap-opening mass expression of 1986lin [^13] Equation 34 of 2014schleicher has a missing factor of $\sqrt{\pi}$ compared to that of 1986lin.. They obtained masses of 6.6 MJup and 1.9 MJup (cf. observed values of 7 MJup and 1.7 MJup) for orbital separations of 5.4 au and 3.4 au.
@@
We then estimate the final clump mass (Equations 30 and 31) by using a realistic temperature profile and not assuming that $Q=1$ (Equation 7). The predicted final clump masses for orbital separations of 5.4 au and 3.4 au (blue line) are 2 MJup and 1.2 MJup, somewhat lower than the values obtained by 2014schleicher. Despite this small discrepancy, the analysis of 2014schleicher is consistent with what we have found, so that we can conclude that if the disc mass is close to their adopted 0.1 M then indeed this disc could make the observed planets by gravitational fragmentation.
Figure 11: Clump masses and final (planet) masses as a function of orbital distance for NN Ser using the values used by 2014schleicher, but assuming the temperature profile of Equation 7 and $r_{\rm out}=7.5$ au.